# On $f$-Divergence Principled Domain Adaptation: An Improved Framework

### v1

@@ 5 Sharper Bounds via Localization
ResNet-50 (He et al., 2016) 68.4$\pm$0.2 96.7$\pm$0.1 99.3$\pm$0.1 68.9$\pm$0.2 62.5$\pm$0.3 60.7$\pm$0.3 76.1
DANN (Ganin et al., 2016) 82.0$\pm$0.4 96.9$\pm$0.2 99.1$\pm$0.1 79.7$\pm$0.4 68.2$\pm$0.4 67.4$\pm$0.5 82.2
MDD (Zhang et al., 2019) 94.5$\pm$0.3 98.4$\pm$0.1 100.0$\pm$.0 93.5$\pm$0.2 74.6$\pm$0.3 72.2$\pm$0.1 88.9
KL (Nguyen et al., 2022) 87.9$\pm$0.4 99.0$\pm$0.2 100.0$\pm$.0 85.6$\pm$0.6 70.1$\pm$1.1 69.3$\pm$0.7 85.3
$f$-DAL (Acuna et al., 2021) 95.4$\pm$0.7 98.8$\pm$0.1 100.0$\pm$.0 93.8$\pm$0.4 74.9$\pm$1.5 74.2$\pm$0.5 89.5
@@
ResNet-50 (He et al., 2016) 34.9 50.0 58.0 37.4 41.9 46.2 38.5 31.2 60.4 53.9 41.2 59.9 46.1
DANN (Ganin et al., 2016) 45.6 59.3 70.1 47.0 58.5 60.9 46.1 43.7 68.5 63.2 51.8 76.8 57.6
MDD (Zhang et al., 2019) 54.9 73.7 77.8 60.0 71.4 71.8 61.2 53.6 78.1 72.5 60.2 82.3 68.1
$f$-DAL (Acuna et al., 2021) 54.7 71.7 77.8 61.0 72.6 72.2 60.8 53.4 80.0 73.3 60.6 83.8 68.5
@@
6 Algorithms and Experimental Results
6.1 Domain Adversarial Learning Algorithm
Adversarial Learning Obejective
In a practical algorithm, the hypothesis space consists of two components: the representation part, denoted as $\mathcal{H}_{\rm rep}=\{h_{\rm rep}:\mathcal{X}\to\mathcal{Z}\}$, where $\mathcal{Z}$ is the representation space (e.g., the hidden output of a neural network), and the classification part, denoted as $\mathcal{H}_{\rm cls}=\{h_{\rm cls}:\mathcal{Z}\to\mathcal{Y}\}$. Therefore, the entire hypothesis space is given by =$\mathcal{H}=\{h_{\rm cls}\circ h_{\rm rep}|h_{\rm rep}\in\mathcal{H}_{\rm rep},h_{\rm cls}\in\mathcal{H}_{\rm cls}\}$. The training objective in $f$-DAL (Acuna et al., 2021) is then formulated as:
Eq. (6) results in an adversarial training strategy. Specifically, the outer optimization spans the entire hypothesis space. Meanwhile, within the inner optimization, given a $h=h_{\rm rep}\circ h_{\rm cls}$, the representation component $h_{\rm rep}$ is shared for $h^{\prime}$. In other words, the optimization is carried out for $h^{\prime}$ in $\mathcal{H}^{\prime}=\mathcal{H}_{\rm cls}\circ h_{\rm rep}\triangleq\{h_{\rm cls}\circ h_{\rm rep}|h_{\rm cls}\in\mathcal{H}_{\rm cls}\}$.
Clearly, as also discussed previously, $\max_{h^{\prime}}\tilde{d}_{\hat{\mu},\hat{\nu}}(h,h^{\prime})\leq\max_{h^{\prime}}\left|\tilde{d}_{\hat{\mu},\hat{\nu}}(h,h^{\prime})\right|$, which presents a clear gap between the theory and algorithms in Acuna et al. (2021). In contrast, this training objective aligns more closely with our $\mathrm{D}_{\phi}^{h,\mathcal{H}}$. Formally, we have the following result.
Proposition 1.
Let $d_{\hat{\mu},\hat{\nu}}(h,h^{\prime})={{\mathbb{E}_{\hat{\nu}}}\left[\hat{\ell}(h,h^{\prime})\right]}-I_{\phi,\hat{\mu}}^{h}(\hat{\ell}\circ h^{\prime})$. Assume ^$\hat{\ell}:\mathcal{X}\to\mathbb{R}$, we have $\max_{h^{\prime}}\tilde{d}_{\hat{\mu},\hat{\nu}}(h,h^{\prime})=\max_{h^{\prime}}d_{\hat{\mu},\hat{\nu}}(h,h^{\prime})=\mathrm{D}_{\phi}^{h,\mathcal{H}^{\prime}}(\hat{\nu}||\hat{\mu})\leq\mathrm{D}_{\phi}(\hat{\nu}||\hat{\mu})$.
Discrepancy Measures
In our algorithms, we mainly focus on three specific discrepancy measures: KL-DD, $\chi^{2}$-DD and the weighted Jeffereys-DD. Specifically, the weighted Jeffereys-DD is $\gamma_{1}\mathrm{D}_{\rm KL}(\hat{\mu}||\hat{\nu})+\gamma_{2}\mathrm{D}_{\rm KL}(\hat{\nu}||\hat{\mu})$, where $\gamma_{1}$ and $\gamma_{2}$ are tunable hyper-parameters. We note that Jeffereys divergence is not explored in Acuna et al. (2021) while it is also an $f$-divergence with $\phi(x)=(x-1)\log{x}$, and advantages of Jeffereys divergence are studied in (Bu et al., 2022; Nguyen et al., 2022; Wang & Mao, 2023a).
6.2 Experiments
@@
Baselines and Implementation Details
The primary baseline for our study is the preceding $f$-DAL (Acuna et al., 2021). Note that, with the exception of Digits, the results reported by Acuna et al. (2021) for $f$-DAL rely on maximum values obtained from their $\chi^{2}$-divergence and weighted Jensen-Shannon divergence across individual sub-tasks. In our comparison, we also include DANN (Ganin et al., 2016) and MDD (Zhang et al., 2019) as they may be interpreted as special cases of $f$-DAL. Furthermore, we compare the results reported by Nguyen et al. (2022) on the Office-31 dataset, although denoted as “KL”, their method incorporates Jeffereys divergence in their algorithms.
Our implementation closely follows Acuna et al. (2021). For Office-31 and Office-Home, we utilize a pretrained ResNet-50 (He et al., 2016) as the backbone network, while both the primary classifier in $h$ and the auxiliary classifier in $h^{\prime}$ consist of two-layer Leaky-ReLU networks. In the case of Digits, we employ LeNet (LeCun et al., 1998) as the backbone network and a two-layer ReLU network with Dropout ($0.5$) for the classifiers. Other hyper-parameter settings and the evaluation protocol remain consistent with Acuna et al. (2021), and the reported results represent average accuracies over $3$ different random seeds.
@@
Tables 1-3 collectively demonstrate the superior performance of our weighted Jeffereys-DD compared to other methods across the three benchmarks. Notably, the most significant improvement over $f$-DAL is observed on Office-Home ($70.2\%$ vs $68.5\%$). Remarkably, this performance even surpasses the combination of $f$-DAL with a sampling-based implicit alignment approach (Jiang et al., 2020) (See Table 6 in Appendix), specifically designed to handle the label shift issues. In addition, unlike findings in Acuna et al. (2021), where $\chi^{2}$ outperforms other methods on nearly all tasks, our use of a tighter variational representation-based discrepancy reveals that $\chi^{2}$ is no longer superior to KL. In fact, our KL-DD slightly outperforms $\chi^{2}$-DD across all three benchmarks.
@@
In the paragraph following Proposition 1, we discuss the observation that optimizing over $t$ may not be necessary. Empirical evidence indicates that setting $t=1$ with hyper-parameter tuning (e.g., through $\eta$) yields satisfactory performance. Now, let’s delve into the selection of $t$ for KL-DD. Instead of using a stochastic gradient-based optimizer for updating $t$, we opt for a quadratic approximation for the optimal $t$, as studied in Birrell et al. (2022). Specifically, we define a Gibbs measure $d\mu^{\prime}=\frac{e^{{\ell}(h,h^{\prime})}d\mu}{{{\mathbb{E}_{\mu}}\left[e^{{\ell}(h,h^{\prime})}\right]}}$, then the optimal $t^{*}\approx 1+\Delta t^{*}$, where $\Delta t^{*}=\frac{{{\mathbb{E}_{\nu}}\left[{\ell}(h,h^{\prime})\right]}-{{\mathbb{E}_{\mu^{\prime}}}\left[\ell(h,h^{\prime})\right]}}{\mathrm{Var}_{\mu^{\prime}}(\ell(h,h^{\prime}))}$. Interested readers can find a detailed derivation of this approximation in Birrell et al. (2022, Appendix B). Substituting $t=1+\Delta t^{*}$, we have the training objective for the approximately optimal KL-DD (OptKL-DD). Table 4 presents an empirical comparison between OptKL-DD and the original KL-DD, where $t$ is simply set to 1. The results indicate that OptKL-DD does not offer any improvement on these benchmarks. Similar observations also hold for $\chi^{2}$, in which case optimal $t$ has an analytic form (see Appendix E), suggesting that using $t=1$, at least for KL and $\chi^{2}$, might be sufficient in practice.
Additional empirical results, including an ablation study on $\eta$ and t-SNE (Van der Maaten & Hinton, 2008) visualizations, can be found in Appendix E.
@@ 7 Other Related Works
Apart from those mentioned in the introduction, various other discrepancy measures are explored in DA theories and algorithms. These include the Wasserstein distance (Flamary et al., 2016; Courty et al., 2017; Shen et al., 2018), Maximum Mean Discrepancy (Huang et al., 2006; Gong et al., 2013), second-order statistics alignment (Sun & Saenko, 2016; Sun et al., 2016), transfer exponents (Hanneke & Kpotufe, 2019; Hanneke et al., 2023), Integral Probability Metrics (Dhouib & Maghsudi, 2022) and so on. Additionally, one of our baseline models Nguyen et al. (2022) diverges from the adversarial training strategy. Instead, they minimize the KL divergence between two isotropic Gaussian distributions (source domain Gaussian and target domain Gaussian) in the representation space. Here, the Gaussian means and variances correspond to the hidden outputs of the representation network. The comparison between their KL-guided framework and our KL-DD is akin to the distinction between VAEs (Kingma & Welling, 2013) and GANs (Goodfellow et al., 2014). For further literature on DA theory, readers are directed to a recent survey by Redko et al. (2020).
@@
The combination of $f$-divergence and adversarial training schemes has been extensively studied in generative models, including $f$-GAN (Nowozin et al., 2016; Song & Ermon, 2020), $\chi^{2}$-GAN (Tao et al., 2018) and others. In the DA context, Wu et al. (2019) introduce a $f$-divergence-based discrepancy measure while still relying on Lemma 2.1 and focusing solely on the Jensen-Shannon case. Additionally, Mansour et al. (2009b) investigates $\alpha$-Rényi divergence for multi-source DA, and Bruns-Smith et al. (2022) provides some intriguing interpretations of $\chi^{2}$-divergence-based generalization bound for covariate shifts.
@@ 8 Conclusion and Future Work
We have presented an improved approach for integrating $f$-divergence into DA theory. Theoretical contributions include novel DA generalization bounds, including fast-rate bounds via localization. On the practical front, the revised $f$-divergence-based discrepancy improves benchmark performance. Several promising future directions emerge from our work. Firstly, beyond its usefulness for local Rademacher complexity, the Rashomon set $\mathcal{H}_{r}$ also relates to another generalization measure, Rashomon ratio (Semenova et al., 2022), which may give an alternative perspective on generalization in DA. Additionally, exploring transfer component-based analysis (Hanneke & Kpotufe, 2019) for tight minimax rates in DA, leveraging a power transformation instead of the affine transformation in $f$-DD, holds promise.
@@ Appendix A Summary of Notations
$\ell(h(x),h^{\prime}(x))$ or $\ell(h,h^{\prime})$ loss between $h$ and $h^{\prime}$
@@
^$\hat{\mathfrak{R}}_{S}(\mathcal{F})$ ${{\mathbb{E}_{\varepsilon_{1:n}}}\left[\sup_{f\in\mathcal{F}}{\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(Z_{i})}\right]}$; Rademacher complexity
@@
$\hat{\ell}$ surrogate loss used in practical algorithm
@@ Appendix B Some Technical Lemmas
Definition B.1 (Rademacher Complexity (Bartlett & Mendelson, 2002)).
@@
The Rademacher complexity based generalization bound is given below.
Lemma B.3 (Mohri et al. (2018, Theorem 3.3)).
@@
The following result is from Agrawal & Horel (2020, Proposition 4.1.), and the corresponding detailed proof is given in Agrawal & Horel (2021, Corollary 6.3.11).
@@
Assume that $\phi$ is twice differentiable on its domain and ϕ′′superscriptitalic-$\phi^{\prime\prime}$ is monotone. Let $\psi(x)=\phi(x+1)$ and let $f:\mathcal{X}\to[a,b]$, then for each $t\in\mathbb{R}$, we have $\inf_{\alpha}{{\mathbb{E}_{X}}\left[\psi^{*}\left(tf(X)+\alpha\right)\right]}\leq\frac{(b-a)^{2}}{2\phi^{\prime\prime}(1)}\cdot t^{2}$.
@@
${\mathbb{E}\left[e^{X}\right]}\leq{\mathbb{E}\left[X\right]}+1+{\mathbb{E}\left[X^{2}B(b)\right]}\leq e^{B(b){\mathbb{E}\left[X^{2}\right]}}.$
@@
Lemma B.6 (Talagrand’s inequality (Boucheron et al., 2005, Theorem 5.4)).

### v2

@@ 6.1 Domain Adversarial Learning Algorithm
ResNet-50 [44] 68.4$\pm$0.2 96.7$\pm$0.1 99.3$\pm$0.1 68.9$\pm$0.2 62.5$\pm$0.3 60.7$\pm$0.3 76.1
DANN [12] 82.0$\pm$0.4 96.9$\pm$0.2 99.1$\pm$0.1 79.7$\pm$0.4 68.2$\pm$0.4 67.4$\pm$0.5 82.2
MDD [9] 94.5$\pm$0.3 98.4$\pm$0.1 100.0$\pm$.0 93.5$\pm$0.2 74.6$\pm$0.3 72.2$\pm$0.1 88.9
KL [45] 87.9$\pm$0.4 99.0$\pm$0.2 100.0$\pm$.0 85.6$\pm$0.6 70.1$\pm$1.1 69.3$\pm$0.7 85.3
$f$-DAL [1] 95.4$\pm$0.7 98.8$\pm$0.1 100.0$\pm$.0 93.8$\pm$0.4 74.9$\pm$1.5 74.2$\pm$0.5 89.5
@@
ResNet-50 [44] 34.9 50.0 58.0 37.4 41.9 46.2 38.5 31.2 60.4 53.9 41.2 59.9 46.1
DANN [12] 45.6 59.3 70.1 47.0 58.5 60.9 46.1 43.7 68.5 63.2 51.8 76.8 57.6
MDD [9] 54.9 73.7 77.8 60.0 71.4 71.8 61.2 53.6 78.1 72.5 60.2 82.3 68.1
$f$-DAL [1] 54.7 71.7 77.8 61.0 72.6 72.2 60.8 53.4 80.0 73.3 60.6 83.8 68.5
@@
Eq. (5) results in an adversarial training strategy. Specifically, the outer optimization spans the entire hypothesis space. Meanwhile, within the inner optimization, given a $h=h_{\rm rep}\circ h_{\rm cls}$, the representation component $h_{\rm rep}$ is shared for $h^{\prime}$. In other words, the optimization is carried out for $h^{\prime}$ in $\mathcal{H}^{\prime}=\mathcal{H}_{\rm cls}\circ h_{\rm rep}\triangleq\{h_{\rm cls}\circ h_{\rm rep}|h_{\rm cls}\in\mathcal{H}_{\rm cls}\}$. The overall training framework of our $f$-DD is illustrated in Figure 2 in Appendix.
Clearly, as also discussed previously, $\max_{h^{\prime}}\tilde{d}_{\hat{\mu},\hat{\nu}}(h,h^{\prime})\leq\max_{h^{\prime}}\left|\tilde{d}_{\hat{\mu},\hat{\nu}}(h,h^{\prime})\right|$, which presents a clear gap between the theory and algorithms in [1]. In contrast, this training objective aligns more closely with our $\mathrm{D}_{\phi}^{h,\mathcal{H}}$. Formally, we have the following result.
Proposition 1.
Let $d_{\hat{\mu},\hat{\nu}}(h,h^{\prime})={{\mathbb{E}_{\hat{\nu}}}\left[\hat{\ell}(h,h^{\prime})\right]}-I_{\phi,\hat{\mu}}^{h}(\hat{\ell}\circ h^{\prime})$. Assume $\mathcal{H}$ is sufficiently large s.t. ^$\hat{\ell}:\mathcal{X}\to\mathbb{R}$, we have $\max_{h^{\prime}}\tilde{d}_{\hat{\mu},\hat{\nu}}(h,h^{\prime})=\max_{h^{\prime}}d_{\hat{\mu},\hat{\nu}}(h,h^{\prime})=\mathrm{D}_{\phi}^{h,\mathcal{H}^{\prime}}(\hat{\nu}||\hat{\mu})\leq\mathrm{D}_{\phi}(\hat{\nu}||\hat{\mu})$.
Furthermore, we highlight that the training objective in our algorithm, namely Eq. (5) with $\tilde{d}_{\hat{\mu},\hat{\nu}}(h,h^{\prime})$ replaced by $d_{\hat{\mu},\hat{\nu}}(h,h^{\prime})$, is de facto motivated by the insights from Section 5. In that section, we demonstrate that when the source domain error is small, the $f$-DD term without the square-root function provides a more accurate reflection of generalization (see, for instance, Remark 5.3). Given that the empirical risk of the source domain is always minimized during training, we expect the final hypothesis to fall within the subset of $\mathcal{H}$ (i.e. Rashomon set with $r_{1}$).
6.2 Experiments
@@
Discrepancy Measures
In our algorithms, we mainly focus on three specific discrepancy measures: KL-DD, $\chi^{2}$-DD and the weighted Jeffereys-DD. Specifically, the weighted Jeffereys-DD is $\gamma_{1}\mathrm{D}_{\rm KL}(\hat{\mu}||\hat{\nu})+\gamma_{2}\mathrm{D}_{\rm KL}(\hat{\nu}||\hat{\mu})$, where $\gamma_{1}$ and $\gamma_{2}$ are tunable hyper-parameters. We note that Jeffereys divergence is not explored in [1] while it is also an $f$-divergence with $\phi(x)=(x-1)\log{x}$, and advantages of Jeffereys divergence are studied in [18, 45, 19].
Baselines and Implementation Details
The primary baseline for our study is the preceding $f$-DAL [1]. Note that, with the exception of Digits, the results reported by [1] for $f$-DAL rely on maximum values obtained from their $\chi^{2}$-divergence and weighted Jensen-Shannon divergence across individual sub-tasks. In our comparison, we also include DANN [12] and MDD [9] as they may be interpreted as special cases of $f$-DAL. Furthermore, we compare the results reported by [45] on the Office-31 dataset, although denoted as “KL”, their method incorporates Jeffereys divergence in their algorithms.
Our implementation closely follows [1]. For Office-31 and Office-Home, we utilize a pretrained ResNet-50 [44] as the backbone network, while both the primary classifier in $h$ and the auxiliary classifier in $h^{\prime}$ consist of two-layer Leaky-ReLU networks. In the case of Digits, we use LeNet [50] as the backbone network and a two-layer ReLU network with Dropout ($0.5$) for the classifiers. Other hyper-parameter settings and the evaluation protocol remain consistent with [1], and the reported results represent average accuracies over $3$ different random seeds.
@@
Tables 1-3 collectively demonstrate the superior performance of our weighted Jeffereys-DD compared to other methods across the three benchmarks. Notably, the most significant improvement over $f$-DAL is observed on Office-Home ($70.2\%$ vs. $68.5\%$). Remarkably, this performance even surpasses the combination of $f$-DAL with a sampling-based implicit alignment approach [51] (See Table 6 in Appendix), specifically designed to handle the label shift issues. In addition, unlike findings in [1], where $\chi^{2}$ outperforms other methods on nearly all tasks, our use of a tighter variational representation-based discrepancy reveals that $\chi^{2}$ is no longer superior to KL. In fact, our KL-DD slightly outperforms $\chi^{2}$-DD in all three benchmarks.
@@
In the paragraph following Proposition 1, we discuss the observation that optimizing over $t$ may not be necessary. Empirical evidence indicates that setting $t=1$ with hyper-parameter tuning (e.g., through $\eta$) obtains satisfactory performance. Now, let’s investigate the selection of $t$ for KL-DD. Instead of using a stochastic gradient-based optimizer for updating $t$, we invoke a quadratic approximation for the optimal $t$, as studied in [23]. Specifically, we define a Gibbs measure $d\mu^{\prime}=\frac{e^{{\ell}(h,h^{\prime})}d\mu}{{{\mathbb{E}_{\mu}}\left[e^{{\ell}(h,h^{\prime})}\right]}}$, then the optimal $t^{*}\approx 1+\Delta t^{*}$, where $\Delta t^{*}=\frac{{{\mathbb{E}_{\nu}}\left[{\ell}(h,h^{\prime})\right]}-{{\mathbb{E}_{\mu^{\prime}}}\left[\ell(h,h^{\prime})\right]}}{\mathrm{Var}_{\mu^{\prime}}(\ell(h,h^{\prime}))}$. Interested readers can find a detailed derivation of this approximation in [23, Appendix B]. Substituting $t=1+\Delta t^{*}$, we have the training objective for approximately optimal KL-DD (OptKL-DD). Table 4 presents an empirical comparison between OptKL-DD and the original KL-DD, where $t$ is simply set to 1. The results indicate that OptKL-DD does not provide any improvement on these benchmarks. Similar observations also hold for $\chi^{2}$, in which case optimal $t$ has an analytic form (see Appendix E), suggesting that using $t=1$, at least for KL and $\chi^{2}$, might be sufficient in practice.
Additional experimental results, including an ablation study on $\eta$, t-SNE [52] visualizations and other comparisons, can be found in Appendix E.
@@ 7 Other Related Works
Apart from those mentioned in the introduction, various other discrepancy measures are explored in DA theories and algorithms. These include the Wasserstein distance [53, 54, 55], Maximum Mean Discrepancy [56, 57], second-order statistics alignment [58, 59], transfer exponents [60, 61], Integral Probability Metrics [62] and so on. Notably, [62] defines a general Integral Probability Metrics (IPMs)-based discrepancy measure. It’s noteworthy that the intersection of the IPMs family and the $f$-divergence family results in the total variation. Consequently, both corresponding discrepancy measures can consider $\mathcal{H}\Delta\mathcal{H}$-divergence as a special case. Additionally, one of our baseline models [45] diverges from the adversarial training strategy. Instead, they directly minimize the KL divergence between two isotropic Gaussian distributions (source domain Gaussian and target domain Gaussian) in the representation space. Here, the Gaussian means and variances correspond to the hidden outputs of the representation network. For further literature on DA theory, readers are directed to a recent survey by [63].
@@
Moreover, the combination of $f$-divergence and adversarial training schemes has been extensively studied in generative models, including $f$-GAN [64, 65], $\chi^{2}$-GAN [66] and others. In the DA context, [67] introduce a $f$-divergence-based discrepancy measure while still relying on Lemma 2.1 and focusing solely on the Jensen-Shannon case. Additionally, [68] investigates $\alpha$-Rényi divergence for multi-source DA, and [69] provides some intriguing interpretations of $\chi^{2}$-divergence-based generalization bound for covariate shifts.
@@ 8 Conclusion and Future Work
In this work, we present an improved approach for integrating $f$-divergence into DA theory. Theoretical contributions include novel DA generalization bounds, including fast-rate bounds via localization. On the practical front, the revised $f$-divergence-based discrepancy improves the benchmark performance. Several promising future directions emerge from our work. Firstly, beyond its usefulness for local Rademacher complexity, the Rashomon set $\mathcal{H}_{r}$ also relates to another generalization measure, Rashomon ratio [37], which may give an alternative perspective on generalization in DA. Additionally, exploring transfer component-based analysis [60] for tight minimax rates in DA, invoking a power transformation instead of the affine transformation in $f$-DD, holds promise.
Acknowledgments and Disclosure of Funding
This work is supported partly by an NSERC Discovery grant. The authors would like to thank Loïc Simon for bringing reference [35] to their attention and for the insightful feedback provided on this work. The authors also thank Fanshuang Kong for the extensive experimental guidance provided throughout this project. Furthermore, the authors are thankful to the anonymous AC and reviewers for their careful reading and valuable suggestions.
@@ Appendix A Summary of Notations
$\ell(h(x),h^{\prime}(x))$ or $\ell(h,h^{\prime})$ loss between the predictions returned by $h$ and $h^{\prime}$
@@
^$\hat{\mathfrak{R}}_{S}(\mathcal{F})$ ${{\mathbb{E}_{\varepsilon_{1:n}}}\left[\sup_{f\in\mathcal{F}}{\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(Z_{i})}\right]}$; (empirical) Rademacher complexity
@@
$\hat{\ell}$ surrogate loss used in practical algorithms
@@ Appendix B Some Technical Lemmas
Definition B.1 (Empirical Rademacher Complexity [34]).
@@
The Rademacher complexity-based generalization bound is given below.
Lemma B.3 ([70, Theorem 3.3]).
@@
The following result is from [21, Proposition 4.1.], and the corresponding detailed proof is given in [22, Corollary 6.3.11].
@@
Assume that $\phi$ is twice differentiable on its domain and $\phi^{\prime\prime}$ is monotone. Let $\psi(x)=\phi(x+1)$ and let $f:\mathcal{X}\to[a,b]$, then for each $t\in\mathbb{R}$, we have $\inf_{\alpha}{{\mathbb{E}_{X}}\left[\psi^{*}\left(tf(X)+\alpha\right)\right]}\leq\frac{(b-a)^{2}}{2\phi^{\prime\prime}(1)}\cdot t^{2}$.
@@
${\mathbb{E}\left[e^{X}\right]}\leq{\mathbb{E}\left[X\right]}+1+{\mathbb{E}\left[X^{2}B(b)\right]}\leq e^{B(b){\mathbb{E}\left[X^{2}\right]}}.$ (6)
@@
Lemma B.6 (Talagrand’s inequality [72, Theorem 5.4]).